OpenStudy (anonymous):

Do anyone know the answer to -3/4m+3=8?

OpenStudy (anonymous):

yep its 20/3

TheSmartOne (thesmartone):

\[-\frac{ 3 }{ 4 }m+3=8\] \[-\frac{ 3 }{ 4 }m=5\] \[-\frac{ 4 }{ 3 }\times -\frac{ 3 }{ 4 }\times m=5 \times -\frac{ 4 }{ 3 }\] \[m=-\frac{ 20 }{3 }\]